. Assume the following for a stock and a call and a put option written on the stock.

EXERCISE PRICE = $20

CURRENT STOCK PRICE = $22

VARIANCE = .25

Standard Deviation = .50

TIME TO EXPIRATION = 4 MONTHS  T = .33

RISK FREE RATE = 3%

Use the Black Scholes procedure to determine the value of the call option and the value of a put.
Answer #1
S = Current Stock Price = 22
t = time until option expiration(years) = 0.3300
X = Option Strike Price = 20
r = risk free rate(annual) = 0.0300
s = standard deviation(annual) = 0.50
N = cumulative standard normal distribution
d1 = {ln (S/K) + (r +s^2/2)t}/s√t
= {ln (22/20) + (0.03 + 0.5^2/2)*0.33}/0.5*√0.33
0.509900
d2 = d1 - s√t
= 0.5099 - 0.5√0.33
0.2227
Using z tables,
N(d1) = 0.6949
N(d2) = 0.5881
C = Call Premium = =SN(d1) - N(d2)Ke^(-rt)
= 22*0.6949 - 0.5881*20e^(-0.03*0.33)
3.6417
N(-d1) = 0.3051
N(-d2) = 0.4119
P = Put Premium = =N(-d2)Ke^(-rt) - SN(-d1)
= 0.4119*20e^(-0.03*0.33) - 22*0.3051
1.4446

Hence, value of call option = $3.64 and value of put option = $1.44